OpenStudy (anonymous):

If x varies directly with y and x = 2.5 when y = 10, find x when y = 16.

OpenStudy (anonymous):

in direct variation x1/y1 = x2/y2 so 2.5 x2 2.5 * 16 --- = ---- ==> ---------- = ? ==> 4 = x2 10 16 10

OpenStudy (radar):

\[x = y/k\] \[2.5=10/k\] \[k=10/2.5=4\]

OpenStudy (radar):

\[x=16/4=4\]

OpenStudy (radar):

Or you could say:\[x=ky\]\[2.5=10k\]\[k=2.5/10=.25\] \[x=.25 (16)=4\] Same answer, but more user friendly.